# Escaping the Verifier: Learning to Reason via Demonstrations

### v1

@@ C.1 Stable & Efficient Learning
While Algorithm 1 provides a practical procedure for alternating updates of the policy ($\theta$) and the critic model ($\phi$) it requires training two reasoning LLMs and thus incurs long, token-intensive rollouts for both. To reduce memory usage and potentially promote generalization via shared representations, we ultimately use the same underlying LLM to role-play as both the critic and the policy. Our ablations (see Appendix D.2) empirically support that using a shared LLM for the critic and the policy improves performance.
@@ C.3 Implementation Stack
We train the SFT and Rationalization baselines using Together AI’s managed fine-tuning service. While we monitor the validation loss during training, we ultimately select the checkpoitn for evaluation based on the best validation reward.
Iteractive Direct Preference Optimization (DPO).
@@ D.1 Model Size Scaling on Countdown
We systematically study the effect of scaling model size on the performance on Countdown. In addition to the main results at 1.5B reported in Section 4, we conducted additional experiments at 3B. The verifiable baseline, RLVR, exhibits a performance regression, dropping from $57.7\%$ at 1.5B to $53.5\%$ at 3B (Figure 10). Similarly, we observe that the performance of RARO also degrades from $54.4\%$ to $49.7\%$ at 3B. Furthermore, as illustrated in Figure 10, after inital improvements, both RLVR and RARO performance actively decreases as trainiing progresses. While we do not have a definitive explanation, we hypothesize that larger models may be more prone to the training-inference log-probability mismatch problem (yao2025offpolicy), leading to degradation when scaling model capacity. These results indicate that RARO does not inherently contribute to the performance plateau; rather, it is a systematic problem that we observe with RLVR as well.
@@ Appendix E Additional Tables & Figures
Algorithm 4 Alternating Reward-Policy Optimization
Models: Reward $r_{\phi}(a,q)$; Policy $\pi_{\theta}(z,a\mid q)$.
1:Initialize $\phi,\theta$
2:for $t=1,\ldots,T$ do
5:  Reward update:
6:  Policy update: KL-regularized reward-maximization with $r_{\phi}(a,q)$ as the reward.
7:end for
8:return $\theta,\phi$
@@
accuracy ($\%$) $\uparrow$ score (0-100) $\uparrow$ win-rate ($\%$) $\uparrow$
@@
accuracy ($\%$) $\uparrow$ score (0-100) $\uparrow$ win-rate ($\%$) $\uparrow$
@@
Figure 15: Full qualitative example for Countdown policy at 1.5B.
DeepMath Input Find a positive integer $n>1$ that is both a sum of two squares ($n=a^{2}+b^{2}$) and a perfect cube ($n=z^{3}$), where $a\neq b$. Determine the non-trivial solutions to the equation $a^{2}+b^{2}=z^{3}$.
DeepMath Policy Response <think> To find a positive integer $n>1$ that is both a sum of two squares and a perfect cube, where $a\neq b$, we start by considering the general form of such numbers. A number $n$ that is both a sum of two squares and a perfect cube can be expressed as $n=z^{3}=(a^{2}+b^{2})$. First, let’s consider the smallest possible values for $z$ and check if they can be expressed as $a^{2}+b^{2}$ with $a\neq b$. For $z=2$, we need $n=2^{3}=8$. We check if 8 can be expressed as $a^{2}+b^{2}$ with $a\neq b$. The possible pairs $(a,b)$ are $(1,2)$ and $(2,1)$, but since $a\neq b$, this pair does not work. We need to check other values. For $z=3$, we need $n=3^{3}=27$. We check if 27 can be expressed as $a^{2}+b^{2}$ with $a\neq b$. The possible pairs $(a,b)$ are $(1,2\sqrt{3})$ and $(2\sqrt{3},1)$, but since $a$ and $b$ must be integers, this pair does not work. We need to check other values. For $z=4$, we need $n=4^{3}=64$. We check if 64 can be expressed as $a^{2}+b^{2}$ with $a\neq b$. The possible pairs $(a,b)$ are $(2,6)$ and $(6,2)$, but since $a\neq b$, this pair does not work. We need to check other values. For $z=5$, we need $n=5^{3}=125$. We check if 125 can be expressed as $a^{2}+b^{2}$ with $a\neq b$. The possible pairs $(a,b)$ are $(3,4)$ and $(4,3)$, and since $a\neq b$, this pair works. Therefore, $125=3^{2}+4^{2}$. Thus, $n=125$ is a solution to the equation $a^{2}+b^{2}=z^{3}$ with $a\neq b$. We can verify: $125=5^{3}=(3^{2}+4^{2})=9+16=25.$ Therefore, the non-trivial solution is $n=125$. The answer is 125. </think> 125
Figure 16: Full qualitative example for DeepMath policy at 1.5B.
@@
Poetry Input a short free-verse poem about the ironic pageantry of war glory turning to dust at sunset Please carefully think about how to create an artistic poem while adhering to the prompt, and please enclose the final version of your poem (along with the title) in <poem></poem> tags
Poetry Policy Response <think> To write a free-verse poem about war glory, I should first consider the themes of disillusionment, the transience of power, and what aspects of war can be attributed to pageantry. The key elements to include in the poem are: Correspondence and memory The fading glory and aftermath The natural cycle that war disrupts Tackling the irony of war and its ephemeral nature These should be woven into a flowing, rhythmic yet contemplative style, reflective of an artist’s perspective on war. </think> <poem> The Decoration (1942) Nestled, red-glazed in a rail, Is the body of Brigadier Sir Thomas Hassall. A small silver clasp begirt With curly bristles like the bristles On a sparrow’s tail, The officialese hangs from himself: He was never slated. Illinois was known to my grandmothers Across the Atlantic who dreamed Of Christmas in seventeen seventy-seven, To which evening after sunset they got up from dinner To put vivid stars on the canvas and skate On one another’s knees. In Delhi we shall bury this hero. The sun tentatively discloses its hidden finger. Below the silent-grandstand of airships, His very gold is dust. </poem>
@@
Figure 17: Full qualitative example for Poetry Writing policy at 1.5B.
@@
Countdown Critic Input You are an expert mathematics critic whose goal is to compare two answers to a math problem and determine which one is correct. Question: <question> Use the four numbers 20, 45, 10, and 23 to create an expression that equals 24. </question> Answer 1: <answer_1> (4523)+(20/10) </answer_1> Answer 2: <answer_2> (20+452)23 </answer_2>
@@
Countdown Critic Response <think> To determine which answer is correct, let’s carefully analyze both solutions. For Answer 1: (4523)+(20/10) Let’s follow these steps: 1. Perform the subtraction inside the first set of parentheses: 4523=2245-23=22 2. Perform the division inside the second set of parentheses: $20/10=2$ 3. Add the results: $22+2=24$ Answer 1 is correct. For Answer 2: (20+452)23 Let’s follow these steps: 1. Perform the multiplication inside the parentheses: $45\times 2=90$ 2. Add the numbers inside the parentheses: $20+90=110$ 3. Subtract 23: 11023=87110-23=87 Answer 2 is incorrect. </think> Answer 1
Figure 18: Qualitative example for Countdown critic at 1.5B.
@@
DeepMath Critic Input You are an expert mathematics critic whose goal is to compare two answers to a math problem and determine which one is correct. Question: <question> Evaluate the double integral $\int\int_{R}x^{2}+y^{2}\,dx\,dy$ where $R$ is the region bounded by the circle $x^{2}+y^{2}=4$ and the lines $y=0$ and $y=x$. </question> Answer 1: <answer_1> 8π5 </answer_1> Answer 2: <answer_2> π </answer_2>
@@
DeepMath Critic Response <think> Let’s evaluate the double integral $\iint_{R}x^{2}+y^{2}\,dx\,dy$ where $R$ is the region bounded by the circle $x^{2}+y^{2}=4$ and the lines $y=0$ and $y=x$. To solve the integral, we can convert to polar coordinates. The circle $x^{2}+y^{2}=4$ has a radius of 2, so in polar coordinates, the region $R$ is described by $0\leq r\leq 2$ and $0\leq\theta\leq\frac{\pi}{4}$ (considering the quadrant where the region is). The integrand in polar coordinates is $x^{2}+y^{2}=r^{2}$. The Jacobian of the transformation from Cartesian to polar coordinates is $r$, so the integral becomes: $\iint_{R}x^{2}+y^{2}\,dx\,dy=\int_{0}^{\frac{\pi}{4}}\int_{0}^{2}r^{2}\cdot r\,dr\,d\theta=\int_{0}^{\frac{\pi}{4}}\int_{0}^{2}r^{3}\,dr\,d\theta.$ First, we integrate with respect to $r$: $\int_{0}^{2}r^{3}\,dr=\left[\frac{r^{4}}{4}\right]_{0}^{2}=\frac{2^{4}}{4}-\frac{0^{4}}{4}=\frac{16}{4}=4.$ Next, we integrate with respect to $\theta$: $\int_{0}^{\frac{\pi}{4}}4\,d\theta=4\left[\theta\right]_{0}^{\dfrac{\pi}{4}}=4\left(\dfrac{\pi}{4}-0\right)=\pi.$ Therefore, the value of the double integral is $\pi$. So, the answer is $\boxed{\pi}$. </think> Answer 2
Figure 19: Qualitative example for DeepMath critic at 1.5B.
@@
Poetry Critic Input You are an expert poetry critic and professor of creative writing. Your task is to compare two poems that were both written in response to the same prompt. Prompt: <prompt> Write a reflective free-verse poem about remembering the dead and how death feels sudden whether in disasters like the Titanic or in ordinary life. </prompt> Poem 1: <poem_1> On Missing Someone in a Disaster </poem_1> Poem 2: <poem_2> The Titanic </poem_2>
@@
Poetry Critic Response <think> To evaluate these poems, we need to consider each dimension to grasp the key elements and compare them effectively. The prompt is to reflect on the experience of remembering the dead and perceiving death in contrast to major disasters like the Titanic. This prompts us to focus on both human and naturalistic death, the impact of memory, and the honor and acceptance of grief. Both poems satisfy the prompt by offering reflections on the human condition related to death. They both focus on encounters with those who were absent in disasters and how death intrudes on life. However, Poem 1 uses a speculative metaphor (not “The Titanic,” but an abstractus timeframe-filled poem) to avoid repetition of a specific disaster. Poem 2, on the other hand, employs detailed imagery and simile relating directly to the well-known disaster. Both require the same thoughtful consideration of the prompt’s themes. The differences in focus and tone help to distinguish these views. While a survivalist approach might imply a romantic view of hardiness and mortality, this perspective doesn’t rule out emotional reflection. Both poems visit the liminal space of collective memory and societal assumptions about mourning, driven by personal memory and regret. In terms of literary merit and craftsmanship, both poems employ powerful strategies: evocative language, specific details from natural or social settings (both following the prompt well). However, the execution and spatial division among lines in Poem 2 yields a more static and inventive feeling, without many explicit sensory references or rhyme schemes. Poem 1 then suffers from non-explicit time markers and symbols to depict the changing circumstances of disaster and loss. Unlike Poem 2, which doesn’t explicitly use tragedy but rather deals subjectively with grief and loss, Poem 1’s merging of analogies to the tragedy and reality illustrates the suspension and celebration of the poem’s thorough reflection on these serious ideas. </think> Final Judgment: Poem 2
Figure 20: Qualitative example for Poetry critic at 1.5B.

### v3

@@ C.1 Stable & Efficient Learning
While Section 3.3 provides a practical procedure for alternating updates of the policy ($\theta$) and the critic model ($\phi$), it requires training two reasoning LLMs and thus incurs long, token-intensive rollouts for both. To reduce memory usage and potentially promote generalization via shared representations, we ultimately use the same underlying LLM to role-play as both the critic and the policy. Our ablations (see Appendix D.2) empirically support that using a shared LLM for the critic and the policy improves performance.
@@ C.3 Implementation Stack
We train the SFT and Rationalization baselines using Together AI’s managed fine-tuning service. While we monitor the validation loss during training, we ultimately select the checkpoint for evaluation based on the best validation reward.
Iterative Direct Preference Optimization (DPO).
@@ D.1 Model Size Scaling on Countdown
We systematically study the effect of scaling model size on the performance on Countdown. In addition to the main results at 1.5B reported in Section 5, we conducted additional experiments at 3B. The verifiable baseline, RLVR, exhibits a performance regression, dropping from $57.7\%$ at 1.5B to $53.5\%$ at 3B (Figure 11). Similarly, we observe that the performance of RARO also degrades from $54.4\%$ to $49.7\%$ at 3B. Furthermore, as illustrated in Figure 11, after initial improvements, both RLVR and RARO performance actively decreases as training progresses. While we do not have a definitive explanation, we hypothesize that larger models may be more prone to the training-inference log-probability mismatch problem (yao2025offpolicy), leading to degradation when scaling model capacity. These results indicate that RARO does not inherently contribute to the performance plateau; rather, it is a systematic problem that we observe with RLVR as well.
@@ Appendix E Additional Tables & Figures
accuracy ($\%$) $\uparrow$ score (0-100) $\uparrow$
win-rate ($\%$) $\uparrow$
(vs. expert)
@@
accuracy ($\%$) $\uparrow$ score (0-100) $\uparrow$
win-rate ($\%$) $\uparrow$
(vs. expert)
@@
Figure 16: Full qualitative example for the RARO policy for Countdown at 1.5B.
DeepMath Find the equation of the normal line to the curve $(x^{2}+y^{2})^{3/2}=2xy$ at the point $\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$.
RARO Policy <think> To find the equation of the normal line to the curve $(x^{2}+y^{2})^{3/2}=2xy$ at the point $\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$, we need to follow these steps: 1. Differentiate the curve implicitly to find the slope of the tangent line at the given point. 2. Use the slope of the tangent line to find the slope of the normal line. 3. Write the equation of the normal line using the point-slope form. Let’s start with implicit differentiation. Given the equation: $(x^{2}+y^{2})^{3/2}=2xy$ Differentiate both sides with respect to $x$: $\frac{d}{dx}\left((x^{2}+y^{2})^{3/2}\right)=\frac{d}{dx}(2xy)$ Using the chain rule on the left side: $\frac{3}{2}(x^{2}+y^{2})^{1/2}\cdot\frac{d}{dx}(x^{2}+y^{2})=2y+2x\frac{dy}{dx}$ Now, differentiate $x^{2}+y^{2}$ with respect to $x$: $\frac{3}{2}(x^{2}+y^{2})^{1/2}(2x+2y\frac{dy}{dx})=2y+2x\frac{dy}{dx}$ Simplify the left side: $3(x^{2}+y^{2})^{1/2}(x+y\frac{dy}{dx})=2y+2x\frac{dy}{dx}$ Distribute and collect terms involving $\frac{dy}{dx}$ on one side: $3(x^{2}+y^{2})^{1/2}x+3(x^{2}+y^{2})^{1/2}y\frac{dy}{dx}=2y+2x\frac{dy}{dx}$ $3(x^{2}+y^{2})^{1/2}y\frac{dy}{dx}-2x\frac{dy}{dx}=2y-3(x^{2}+y^{2})^{1/2}x$ Factor out $\frac{dy}{dx}$ on the left side: $\frac{dy}{dx}\left(3(x^{2}+y^{2})^{1/2}y-2x\right)=2y-3(x^{2}+y^{2})^{1/2}x$ Solve for $\frac{dy}{dx}$: $\frac{dy}{dx}=\frac{2y-3(x^{2}+y^{2})^{1/2}x}{3(x^{2}+y^{2})^{1/2}y-2x}$ Next, we need to find the slope of the tangent line at the point $\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$: Substitute $x=\frac{1}{\sqrt{2}}$ and $y=\frac{1}{\sqrt{2}}$ into the derivative: $\frac{dy}{dx}=\frac{2\left(\frac{1}{\sqrt{2}}\right)-3\left(\left(\left(\frac{1}{\sqrt{2}}\right)^{2}+\left(\frac{1}{\sqrt{2}}\right)^{2}\right)\right)^{1/2}\left(\frac{1}{\sqrt{2}}\right)}{3\left(\left(\left(\frac{1}{\sqrt{2}}\right)^{2}+\left(\frac{1}{\sqrt{2}}\right)^{2}\right)\right)^{1/2}\left(\frac{1}{\sqrt{2}}\right)-2\left(\frac{1}{\sqrt{2}}\right)}$ Simplify the terms inside: $x^{2}+y^{2}=\left(\frac{1}{\sqrt{2}}\right)^{2}+\left(\frac{1}{\sqrt{2}}\right)^{2}=\frac{1}{2}+\frac{1}{2}=1$ So: $\frac{dy}{dx}=\frac{2\left(\frac{1}{\sqrt{2}}\right)-3\left(1\right)^{1/2}\left(\frac{1}{\sqrt{2}}\right)}{3\left(1\right)^{1/2}\left(\frac{1}{\sqrt{2}}\right)-2\left(\frac{1}{\sqrt{2}}\right)}=\frac{\frac{2}{\sqrt{2}}-\frac{3}{\sqrt{2}}}{\frac{3}{\sqrt{2}}-\frac{2}{\sqrt{2}}}=\frac{\frac{2-3}{\sqrt{2}}}{\frac{3-2}{\sqrt{2}}}=\frac{-1}{1}=-1$ The slope of the tangent line at $\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$ is 1-1. The slope of the normal line is the negative reciprocal of 1-1, which is $1$. Now, use the point-slope form of the equation of a line $y-y_{1}=m(x-x_{1})$, where $m$ is the slope and $(x_{1},y_{1})$ is the point: $y-\frac{1}{\sqrt{2}}=1\left(x-\frac{1}{\sqrt{2}}\right)$ Simplify: $y-\frac{1}{\sqrt{2}}=x-\frac{1}{\sqrt{2}}$ $y=x$ Thus, the equation of the normal line is: $\boxed{y=x}$ </think> y=x
Figure 17: Full qualitative example for the RARO policy for DeepMath at 7B.
@@
Poetry Writing a poem about a child’s fragmented memory of a father leaving, told in intimate, sensory free verse Please carefully think about how to create an artistic poem while adhering to the prompt, and please enclose the final version of your poem (along with the title) in <poem></poem> tags
RARO Policy <think> To create such a poem, I’d: 1. Focus on sensory details that might linger in a child’s memory. 2. Use free verse, with unrhymed, natural cadences to convey the fragmented and stream-of-consciousness nature of recollections. 3. Include elements that represent the end of a relationship, like a suitcase, luggage, or a car, without explicitly stating the scenario. 4. Incorporate physical and emotional sensations that capture the essence of the scene. </think> <poem> One Memory of My Father Leaving There was the sound of a suitcase rolling on the hallway carpet; the thud of it being set down in the sunroom and a shadow of it carried away by someone going somewhere. It smelled of leather, or brown soap, or something carried in it -like salt air, or bare shoulders, or a man thinking and trying not to cry. </poem>
@@
Figure 18: Full qualitative example for the RARO policy for Poetry Writing at 7B.
@@
Countdown You are an expert mathematics critic whose goal is to compare two answers to a math problem and determine which one is correct. Question: <question> Use the four numbers 20, 45, 10, and 23 to create an expression that equals 24. </question> Answer 1: <answer_1> (4523)+(20/10) </answer_1> Answer 2: <answer_2> (20+452)23 </answer_2>
@@
RARO Critic <think> To determine which answer is correct, let’s carefully analyze both solutions. For Answer 1: (4523)+(20/10) Let’s follow these steps: 1. Perform the subtraction inside the first set of parentheses: 4523=2245-23=22 2. Perform the division inside the second set of parentheses: $20/10=2$ 3. Add the results: $22+2=24$ Answer 1 is correct. For Answer 2: (20+452)23 Let’s follow these steps: 1. Perform the multiplication inside the parentheses: $45\times 2=90$ 2. Add the numbers inside the parentheses: $20+90=110$ 3. Subtract 23: 11023=87110-23=87 Answer 2 is incorrect. </think> Answer 1
Figure 19: Qualitative example of the RARO critic for Countdown at 1.5B.
@@
DeepMath You are an expert mathematics critic whose goal is to compare two answers to a math problem and determine which one is correct. Question: <question> Evaluate the double integral $\int\int_{R}x^{2}+y^{2}\,dx\,dy$ where $R$ is the region bounded by the circle $x^{2}+y^{2}=4$ and the lines $y=0$ and $y=x$. </question> Answer 1: <answer_1> 8π5 </answer_1> Answer 2: <answer_2> π </answer_2>
@@
RARO Critic <think> Let’s evaluate the double integral $\iint_{R}x^{2}+y^{2}\,dx\,dy$ where $R$ is the region bounded by the circle $x^{2}+y^{2}=4$ and the lines $y=0$ and $y=x$. To solve the integral, we can convert to polar coordinates. The circle $x^{2}+y^{2}=4$ has a radius of 2, so in polar coordinates, the region $R$ is described by $0\leq r\leq 2$ and $0\leq\theta\leq\frac{\pi}{4}$ (considering the quadrant where the region is). The integrand in polar coordinates is $x^{2}+y^{2}=r^{2}$. The Jacobian of the transformation from Cartesian to polar coordinates is $r$, so the integral becomes: $\iint_{R}x^{2}+y^{2}\,dx\,dy=\int_{0}^{\frac{\pi}{4}}\int_{0}^{2}r^{2}\cdot r\,dr\,d\theta=\int_{0}^{\frac{\pi}{4}}\int_{0}^{2}r^{3}\,dr\,d\theta.$ First, we integrate with respect to $r$: $\int_{0}^{2}r^{3}\,dr=\left[\frac{r^{4}}{4}\right]_{0}^{2}=\frac{2^{4}}{4}-\frac{0^{4}}{4}=\frac{16}{4}=4.$ Next, we integrate with respect to $\theta$: $\int_{0}^{\frac{\pi}{4}}4\,d\theta=4\left[\theta\right]_{0}^{\dfrac{\pi}{4}}=4\left(\dfrac{\pi}{4}-0\right)=\pi.$ Therefore, the value of the double integral is $\pi$. So, the answer is $\boxed{\pi}$. </think> Answer 2
Figure 20: Qualitative example of the RARO critic for DeepMath at 7B.
@@
Poetry Writing You are an expert poetry critic and professor of creative writing. Your task is to compare two poems that were both written in response to the same prompt. Prompt: <prompt> Write a reflective free-verse poem about remembering the dead and how death feels sudden whether in disasters like the Titanic or in ordinary life. </prompt> Poem 1: <poem_1> On Missing Someone in a Disaster </poem_1> Poem 2: <poem_2> The Titanic </poem_2>
@@
RARO Critic <think> To evaluate these poems, we need to consider each dimension to grasp the key elements and compare them effectively. The prompt is to reflect on the experience of remembering the dead and perceiving death in contrast to major disasters like the Titanic. This prompts us to focus on both human and naturalistic death, the impact of memory, and the honor and acceptance of grief. Both poems satisfy the prompt by offering reflections on the human condition related to death. They both focus on encounters with those who were absent in disasters and how death intrudes on life. However, Poem 1 uses a speculative metaphor (not “The Titanic,” but an abstractus timeframe-filled poem) to avoid repetition of a specific disaster. Poem 2, on the other hand, employs detailed imagery and simile relating directly to the well-known disaster. Both require the same thoughtful consideration of the prompt’s themes. The differences in focus and tone help to distinguish these views. While a survivalist approach might imply a romantic view of hardiness and mortality, this perspective doesn’t rule out emotional reflection. Both poems visit the liminal space of collective memory and societal assumptions about mourning, driven by personal memory and regret. In terms of literary merit and craftsmanship, both poems employ powerful strategies: evocative language, specific details from natural or social settings (both following the prompt well). However, the execution and spatial division among lines in Poem 2 yields a more static and inventive feeling, without many explicit sensory references or rhyme schemes. Poem 1 then suffers from non-explicit time markers and symbols to depict the changing circumstances of disaster and loss. Unlike Poem 2, which doesn’t explicitly use tragedy but rather deals subjectively with grief and loss, Poem 1’s merging of analogies to the tragedy and reality illustrates the suspension and celebration of the poem’s thorough reflection on these serious ideas. </think> Final Judgment: Poem 2
Figure 21: Qualitative example of the RARO critic for Poetry Writing at 7B.